OpenStudy (j3nny0rtiz):

please help!!! A train travels 89 kilometers in 3 hours, and then 81 kilometers in 2 hours. What is its average speed? km/hr

OpenStudy (agent0smith):

Total distance divided by total time.

OpenStudy (j3nny0rtiz):

11?

OpenStudy (agent0smith):

Show work.

OpenStudy (nurali):

(dist = speed * time)

OpenStudy (nurali):

d = 89*3 + 81*2 d = _______ km

OpenStudy (nurali):

reply @j3nny0rtiz

OpenStudy (j3nny0rtiz):

i got 429 km

OpenStudy (nurali):

Correct now Find the total time t=3+2 t=_____hrs

OpenStudy (j3nny0rtiz):

5 hrs

OpenStudy (agent0smith):

No @Nurali you misinterpreted the question. They never gave a speed.

OpenStudy (agent0smith):

Add up the given distances. Add up the given times. Divide as i said.

OpenStudy (j3nny0rtiz):

i got 105.2

OpenStudy (agent0smith):

Like I said earlier, show work

OpenStudy (j3nny0rtiz):

i just just the calculator all i need its to know if my answer its correct, i don't need to show work, its for a test.

OpenStudy (agent0smith):

I'm saying show me your work.

OpenStudy (j3nny0rtiz):